A surgical technique is performed on seven patients. You are told there is a 70% chance of success. Find the probability that the surgery is successful for
a.) exactly five patients
b.) at least five patients
c.) less than five patients

Answer by edjones(8007)

Let x=.7 and y=.3
(x+y)^7=x^7+7x^6y+21x^5y^2+...
.
a)21x^5y^2=.3177
.
b)x^7+7x^6y+21x^5y^2=.6471
.
c) 1-.6471=.3529
.
Ed